please tell me various methods to check onto function

Grade:12

1 Answers

Ramesh V
14 years ago

A function f from A to B is called onto if for all b in B there is an a in A such that f (a) = b.All elements in B are used.

Such functions are referred to as surjective.

An example is shown below:

When working in the coordinate plane, the sets A  and B  become the Real numbers, stated as f: R--->R.

This means that for every element in the codomain there exists an element in the domain which maps to it. The method used to establish this property depends on how the function is given and its properties.

Let y be any element of the codomain and x an element of the domain. We solve the equation y = f(x) for x. This gives us a possible candidate for a domain element. We prove it is a suitable domain element by substituting this value into the function.                                                                 As an example, let's prove that f:R-->R defined by f(x) = 3x-4 is onto, where R denotes the real numbers. We let y be a typical element of the codomain and set up the equation y =f(x). then, y = 3x-4 and solving for x we get x =(y+4)/3. Since y is a real number,                                                then (y+4)/3 is a real number and f((y+4)/3) = 3(y+4)/3 -4 = y.

---

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch